OpenStudy (anonymous):

y''+9y=g(t) y(0)=2 y'(0)=-3 H(s) for the system and the impulse response function h(t) and give a formula for the solution to the IVP. I have no idea what to do no examples like this worked in class :/

OpenStudy (anonymous):

these are solutions to the problem... http://prntscr.com/1kkisc

OpenStudy (anonymous):

@Loser66 you know how to work these?

OpenStudy (loser66):

nope, I don't. My prof skip this part, just wrote out the formula as you did. He talked about impulsion in 1 minute

OpenStudy (anonymous):

lol I don't think my professor talked about them at all. He was asking us in class about solving some problems the day after we learned them and was no one was answering questions and he was like "c'mon guys we did these yesterday....like two problems..." lol as if we are supposed to know how to work them after 10 min of exposure to them.

OpenStudy (anonymous):

that's fine im meeting up with a class mate tomorrow and hopefully we can figure it out. thanks for taking a look though

OpenStudy (loser66):

ha, yours is new to me, hehe..

OpenStudy (ash2326):

I can help you find h(s) and h(t). Other parts I don't know

OpenStudy (ash2326):

@rperez36

OpenStudy (loser66):

just post @ash2326 we need it.

OpenStudy (anonymous):

My apologies. This is my husband's post. He just stepped in the shower right quick, but he's getting out. Please help with what you can, I know he would appreciate the help.

OpenStudy (ash2326):

We have \[y'' (t)+9y=g(t)\] Let's take laplace transform of this we know Laplace transform of a derivative of function is s \times the Laplace of a function \[L(f(t))=F(s)\] \[L(f'(t))=s\times F(s)-f(0)\] \[L(f''(t))=s^2\times F(s)-sf(0)-f'(0)\] do you get this part?

OpenStudy (anonymous):

im back @ash2326 and yes I get that part

OpenStudy (ash2326):

just a min

OpenStudy (ash2326):

the answer you have given is that correct? I mean is that given in textbook

OpenStudy (anonymous):

yeah that is what is given

OpenStudy (ash2326):

It seems that it's wrong, it'll be valid on if the initial conditions are zero, y(0)=0 and y'(0)=0 but here we don't have this

OpenStudy (ash2326):

*only

OpenStudy (ash2326):

I'll be back in 10 minutes, could you confirm the question and solution meanwhile?

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

http://prntscr.com/1kksex

OpenStudy (ash2326):

ok, so the question seem correct.

OpenStudy (ash2326):

if y(0)=0 and y'(0)=0 then \[L(y''+9y)=L(g(t))\] \[s^2 Y(s)+9Y(s)=G(s)\] Transfer function is defined as \[\frac{G(s)}{Y(s)}=H(s)\] can you find this from here?

OpenStudy (anonymous):

so h(s)=s2+9?

OpenStudy (ash2326):

oops it's \[\frac{(Y(s))}{(G(s))}=H(s)\]

OpenStudy (anonymous):

I see it says to the ^-1 where is that from?

OpenStudy (ash2326):

Now you try, you'll get

OpenStudy (ash2326):

@rperez36 do you get this? sorry I made a mistake

OpenStudy (anonymous):

it all makes sense except that its I don't see how its \[(s^2+9)^{-1}\]

OpenStudy (anonymous):

as the solution manual states

OpenStudy (anonymous):

s^2+9 ok but to the ^-1 ????

OpenStudy (anonymous):

oh cause your mistake?

OpenStudy (ash2326):

yes, try with the changed formula

OpenStudy (anonymous):

ok so I just take inverse laplace

OpenStudy (ash2326):

yes, now take inverse laplace transform of H(s)

OpenStudy (anonymous):

yes I got what soln says what about for the other solutions?

OpenStudy (ash2326):

those two other parts, I'm sorry but I don' know :(

OpenStudy (anonymous):

ah ok well thank you for helping with what you did. hopefully these wont be on Mondays test

OpenStudy (ash2326):

ha ha :) I hope so